OpenStudy (anonymous):

3Fe (g) + 4H2O (g) --> Fe3O4 (g) + 4H2 (g) When 12.0 mol H2O is mized with 6.0 mol Fe, which is the limiting reactant? What mass of grams in black iron oxide is produced? What's the total mass of Fe needed to form black iron oxide? What is the total mass of water needed to form black oxide? What mass of grams in excess reactant remains when the reaction is completed?

OpenStudy (aaronq):

First off, when you deal with anything in chemical reactions you deal in terms of moles, n. when you wanna find out how many moles something produces you build relationship using the coefficients, like this: \[\frac{ n _{Fe} }{ 3 }=\frac{ n _{H _{2}O} }{ 4}\] for the first: 12 mol H2O/4 = 3 mol and 6 mol Fe/3 = 2 moles therefore the limiting reactant is Fe, since there is less of it available for the rest use the moles of the limiting reactant, since you can't make more than that: \[\frac{ n _{Fe} }{ 3 }=\frac{ n _{Fe _{3}O _{4}} }{ 1 } \rightarrow \frac{ 2mol }{ 3 }=\frac{ n _{Fe _{3}O _{4}} }{ 1 }\]